Apple-Space, Inc. year-end balance sheets are as follows:

At December 31                                                2012                       2011                       2010

Assets

Cash                                                                      33,458                   41,104                   41,570

A/R, Net                                                               96,971                   71,234                   54,318

Merchandise Inventory                                 125,655                 90,440                   58,433

Prepaid Expenses                                            11,104                   10,685                   4,665

Plant Assets, net                                              311,468                 285,378                 260,914

Total Assets                                                       578,656                 498,841                 419,900

Liabilities and Equity

Accounts Payable                                             148,408                 85,990                   54,873

Long-term notes payable secured by

Mortgages on plant assets                           106,612                 114,733                 92,798

Common Stock, $10 per value                    162,500                 162,500                 162,500

Retained Earnings                                            161,136                 135,618                 109,729

Total Liabilities and equity                           578,656                 498,841                 419,900                

Apple-Space Inc. income statements for the years ended December 31, 2012 and 2011, follow

For Year Ended December 31                      2012                       2011

Sales                                                                      752,253                 593,621

COGS                                                                    458,874                 385,854

Other Operating Expenses                           233,198                 150,186

Interest Expense                                              12,788                   13,653

Income Taxes                                                    9,779                     8,904    

Total Costs and Expenses                             714,639                 558,597                                

Net Income                                                        37,614                   35,024

Earnings/Share                                                 2.31                        2.16       

Compute the following:

  1. Debt-to-equity Ratio

Choose Numerator                 /              Choose Denominator                     =             Debt Ratio %

2012       ___________________          __________________                                   _________

2011       ___________________           ___________________                                 _________

  1. Times interest earned

Choose Numerator /      Choose Denominator                     =             Debt Ratio %

2012       ___________________          __________________                                   _________

2011       ___________________           ___________________                                 _________

(Ref 10-11)
Answer #1

Answer:

Debt-to-equity Ratio = Total liabilities / Shareholders’ Equity

2012:

Total liabilities =148,408 + 106,612 = 255,020

Stockholders' equity = Common Stock + Retained Earnings

= 162,500 + 161,136 = 323,636

2011:

Total debt = 85,990 + 114,733 = 200,723  

Stockholders' equity = Common Stock + Retained Earnings

= 162,500 +135,618 = 298,118

Times interest earned = Income before Interest and Taxes or EBIT / Interest Expense

2012:

Income before Interest and Taxes = Sales - COGS - Other Operating Expenses = 752,253 - 458,874 - 233,198 = 60,181

Interest expenses =12,788

2011:

Income before Interest and Taxes =593,621 - 385,854 - 150,186 = 57,581

Interest expenses = 13,653

Debt-to-equity Ratio Total Liabilities/ Shareholders Equity Choose Numerator 255,020 200,723 |/ | Choose Denominator- 323636